OpenStudy (anonymous):

The drawing shows two displacements, 3.0 km due north and 4.0 km due east, and the corresponding resultant vector R. What is the angle θ that R makes with respect to due north?

OpenStudy (anonymous):

Do you do cos 3/5?

OpenStudy (asnaseer):

 In this diagram we can find \(\theta\) from:\[\tan(\theta)=\frac{4}{3}\]